OpenStudy (anonymous):

I'LL Give A Medal! Please Help!! Solve. two x plus four equals eight x equals two x equals four x equals six x equals eight

OpenStudy (anonymous):

Wait wait uh its two cause two times two is for plus four would equal eight

OpenStudy (anonymous):

2x + 4 = 8 We want to isolate the x on the left side, so we have to get rid of that 2 and 4. Which one do we want to eliminate first?

OpenStudy (anonymous):

x is equal to Two 2x= 8 - 4 2x=4 Divide both side 2 then you will get 2 :

OpenStudy (anonymous):

Einstein...the idea is to work through the problems with people step by step, not just give answers.

OpenStudy (cggurumanjunath):